OpenStudy (anonymous):

graph the system of equations. How many solutions exist for this system of equations.

OpenStudy (jdoe0001):

so, where are they?

OpenStudy (anonymous):

me

OpenStudy (jdoe0001):

you said, "graph THE system of equations", which one?

OpenStudy (anonymous):

this y=3x-9 and y=-1/3x+3

OpenStudy (zzr0ck3r):

they are orthogonal lines, so how many times do they cross?

OpenStudy (zzr0ck3r):

orthogonal = perpendicular

OpenStudy (anonymous):

? i don't get how do graph them

OpenStudy (jdoe0001):

pick any two points for each, and plot that

OpenStudy (zzr0ck3r):

do you know how to find the y intercept and the x intercept?

OpenStudy (anonymous):

that is what i have to choose from a.) The system has one solution. b.) There are no solutions to the system. c.) There are an infinite number of solutions

OpenStudy (anonymous):

I am really bad at math

OpenStudy (jdoe0001):

so, just pick any two points for one, plot them on the cartesian plane, do the same for the other, they're both a line, so, all you need is 2 points for each

OpenStudy (anonymous):

so the right one would be A i think

OpenStudy (jdoe0001):

the right one... as opposed to the left one? did you graph them?

OpenStudy (anonymous):

yes but i don't get it i think i did it wrong

OpenStudy (anonymous):

I mean a.) The system has one solution.

OpenStudy (jdoe0001):

well, when you have 2 functions, and "x" and "y" in both usually means, that the "x" and "y" above and below have the same value, or are the same if x = 3 above x= 3 below too a "solution" graphically occurrs, when "x" and "y" have the same value or are the same for both graphics/lines, that is, WHERE they intersect, at that point, "x" and "y" are the same for both so, take a peek at your plotted graph, do they touch each other? more than once? not at all?

OpenStudy (anonymous):

I think they touch once

OpenStudy (jdoe0001):

where? at what coordinates? (x,y)

OpenStudy (anonymous):

x

OpenStudy (jdoe0001):

and what's the value of "x" there?

OpenStudy (anonymous):

3

OpenStudy (jdoe0001):

and "y"?

OpenStudy (anonymous):

3

OpenStudy (jdoe0001):

meaning you haven't done the plotting :/ try plotting them here -> http://www.fooplot.com/

OpenStudy (jdoe0001):

see if they intersect, once, more than once, not at all

OpenStudy (anonymous):

i don't know how to plot the Y=3x-9 y=-1/3x+3

OpenStudy (jdoe0001):

you give a value to "x", say 10, so y will be = 3(10)-9 = 30-9 = 21

OpenStudy (jdoe0001):

"x" is INDEPENDENT to get any value pretty much, "y" is DEPENDENT on whatever "x" gets for its value

OpenStudy (anonymous):

i still don't think i am doing it right on the fooplot

OpenStudy (jdoe0001):

at fooplot all you do is, type it in

OpenStudy (anonymous):

the whole thing

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

they touch almost on 2

OpenStudy (anonymous):

and they only touch once

OpenStudy (anonymous):

thank you :)